Solve by elimination: 3x=2y+7 5x-2y=13

Answer by stanbon(75887)

Solve by elimination:
3x=2y+7
5x-2y=13
----
Rearrange::
3x - 2y = 7
5x - 2y = 13
--------------------
Subtract and solve for "x":
2x = 6
x = 3
-----
Solve for "y"::
3x - 2y = 7
9 - 2y = 7
2y = 2
y = 1
-------------
Cheers,
Stan H.